OpenStudy (anonymous):

What is the slope intercept form of the equation of the line that passes through the point (5,-4) and has a slope of 3/4?

pooja195 (pooja195):

first we need to set it up like this \[\huge~\rm~y-y_1=m(x-x_1)\] whereas m=slope (x1,y1)

OpenStudy (anonymous):

How do you insert the variables into the equation @pooja195

pooja195 (pooja195):

just type it out y-y1=m(x-x1)

OpenStudy (anonymous):

The answer is ZZZZZZZZZZZZZZZZZZZZZZZ

OpenStudy (anonymous):

This question is boring

OpenStudy (anonymous):

I know i really dont want to do it @iamtheultimatehermes

OpenStudy (anonymous):

Lol it is really easy would you like the answer

OpenStudy (anonymous):

sure @iamtheultimatehermes lol

pooja195 (pooja195):

-_- WOW

OpenStudy (anonymous):

wow @iamtheultimatehermes

OpenStudy (anonymous):

Okay may I see your answer options because I have the answer I just want to make sure I am right

OpenStudy (anonymous):

y=-3/4x-1/4 y=3/4x+1/4 y=3/4x+31/4 y=3/4x+31/4 @iamtheultimatehermes

pooja195 (pooja195):

aww he got busy and left :c

pooja195 (pooja195):

Still need help?

OpenStudy (anonymous):

yeah @pooja195

pooja195 (pooja195):

@Librarian wants to do this problem so ill let him do it he loves these ^-^

OpenStudy (anonymous):

okay @pooja195

pooja195 (pooja195):

nvm ok first we need to set it up like this \[\huge~\rm~y-y_1=m(x-x_1)\] whereas m=slope (x1,y1)

OpenStudy (librarian):

@pooja195 you've got this. - Libby

pooja195 (pooja195):

We have this ill plug the info in for you nvm ok \[\huge~\rm~y+4=3/4(x-5)\]

pooja195 (pooja195):

distribute 3/4(x−5)

OpenStudy (anonymous):

Oh okay thanks!! @pooja195

pooja195 (pooja195):

we arent done :)

pooja195 (pooja195):

Did you finish the distributing part?

OpenStudy (anonymous):

yep @pooja195

pooja195 (pooja195):

What did you get? .-.

OpenStudy (anonymous):

3/4x-3 3/4 @pooja195

pooja195 (pooja195):

y+4=3/4x-15/4 i dsitributed whats -15/4-4?

OpenStudy (anonymous):

-7 3/4

pooja195 (pooja195):

y=3/4x+−31/4 thats what u should ahve gotten

OpenStudy (anonymous):

Thank you!! @pooja195

pooja195 (pooja195):

:)